Grade 12th passMagical Mathematics[Interesting Approach]

Is 2log(-5) defined ?.explain?

Gowher Maqbool
11 Years agoGrade 12th pass

3 Answers

Vijay Mukati
11 Years ago
Dear Student, By the defination of the lograthmic, the log function is not defined for the negative values of input variable. i.e. the domain for the log function is the non-negative real numbers. Thanks.
Aakash
11 Years ago
2\log (-5) is not defined as input is -ve
But  \log ((-5)^{2}) is defined as input is +ve
 
And also  2\log (-5) \neq \log ((-5)^{2}) = \log (5^{2})) = 2\log(5)
manmath
10 Years ago
By the defination of the lograthmic, the log function is not defined for the negative values of input variable. i.e. the domain for the log function is the non-negative real numbers. Thanks.